Sarah is 3 times Nick's age. If Sarah is 28 years older than Nick. Find their ages.
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Sarah is 3 times Nick's age. If Sarah is 28 years older than Nick.
Find their ages.
----
S = 3*N
S = N+28
------
Equation:
3N = N+28
2N = 28
Nick = 14 yrs
-----
Sarah = 3*14 = 42 yrs
-------------
Cheers,
Stan H.
---------------

Answer by ikleyn(52781)

.
Let x be the Nick's age.


Then the Sarah age is 3x, according to the condition.


Also, the condition says that 


    3x - x = 28,


which implies  2x = 28  and  x = 14.


Answer.  Nick is 14 years old.  Sarah is 3*14 = 42 years old.

Solved.